The population of the Northgoat neighborhood has been growing steadily since 1990. In 1995, the population was 48000 people. In 2004, it was 58800 people Find an equation in the form of y=mx+b, where x is the number of years past 1990 and y is the population of Northgoat.
I have tried this many ways: (1995,48000) (2004, 58800) getting a slope = 1200
then: y-48000=1200(x-1995), I came up with y=-10800x+102000
HELP, I am not getting this.

Answer by rothauserc(4718)

The key to the x values is "x is the number of years past 1990", therefore the two pairs of data points are (5, 48000) and (14, 58000).
The slope(m) is (58000 - 48000) / (14 - 5) = 10000 / 9
pick the point (5, 48000) and we solve the following for b (y intercept)
standard for for a line is y = mx + b
48000 = (10000/9)5 + b
multiply by 5
48000 = (50000/9) + b
now multiply both sides of = by 9
432000 = 50000 + 9b
9b = 382000
b = 42444.4
the equation is y = (10000/9)x + 42444.4
now we test it with the second point
y = (10000/9)14 + 42444.4 = 57999.9 is approx 58000
the test checks out :-)